[ A plane flies 3600 miles from San Francisco to Tampa Bay in 8 hours, with a tailwind all the way. The return trip on the same route, now with a headwind, takes 9 hours. Assuming both speeds remain constant, find the speed of the plane and the speed of the wind. ]
Answer by TimothyLamb(4379)

s = d / t
---
outbound:
s + w = 3600/8
s + w = 450
---
return:
s - w = 3600/9
s - w = 400
s = 400 + w
---
s + w = 450
400 + w + w = 450
2w = 50
---
w = 25 mph
s = 425 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php